[figure: 9.7 cm]

FERROMAGNETISM IN DILUTED MAGNETIC SEMICONDUCTORS AT LOW CARRIER DENSITY

V. K. Dugaev^{a}, V. I. Litvinov^{b}, J. Barnaś^{c}, A. H. Slobodskyy^{d},W. Dobrowolski^{e}, M. Vieira^{a} aISEL-DEETC, Rua Cons. Emidio Navarro, 1949-014, Lisbon, Portugal bWaveBand Corporation, 375 Van Ness Ave, Torrance, CA 90501, USA

cDept. of Physics, Adam Mickiewicz University, Umultowska 85, 61-614 Poznań, and Institute of Molecular Physics PAS, Smoluchowskiego 17, 60-179 Poznań, Poland dPhysikalisches Institut EP3, Universität Würzburg, 97074 Würzburg, Germany

eInstitute of Physics PAS, al. Lotników 32/46, 02-668 Warsaw, Poland

We will present a field-theoretical approach to ferromagnetism in semiconductors. In this approach disorder is modeled by a random field of impurity concentration. The theory accounts for interacting magnetization fluctuations and disorder corrections. In the mean-field approximation we found that both disorder and interaction of magnetization fluctuations result in suppression of ferromagnetism. We also propose a new mechanism of interaction between magnetic impurities in III-V semiconductors at low carrier concentrations. The mechanism is based on spin dependent hybridization of the acceptor state with the valence band, which is controlled by the magnetic moment of the d-electrons. The model gives ferromagnetic interaction that does not depend on the carrier concentration. The characteristic length of the interaction is determined by the activation energy of the acceptor level. This mechanism can explain ferromagnetism at low density of carriers. In real GaMnAs ferromagnetic semiconductors, both RKKY interaction via free carriers and the proposed mechanism may be equally relevant.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : oral

Corresponding author : V. K. Dugaev

Address for correspondence : Dr. Vitalii Dugaev Department of Electronics and Communications Instituto Superior de Engenharia de Lisboa Rua Conselheiro Emidio Navarro 1949-014 Lisbon, Portugal

Email address : vdugaev@deetc.isel.ipl.pt